A well-known brokerage firm executive claimed that 70% of investors are currently confident of meeting their investment goals. An XYZ Investor Optimism Survey, conducted over a two week period, found that in a sample of 200 people, 71% of them said they are confident of meeting their goals. Test the claim that the proportion of people who are confident is larger than 70% at the 0.05 significance level.

The p-value is: (to 4 decimals)

The null and alternative hypothesis would be:

H0:μ≥0.7
H1:μ<0.7  

H0:p=0.7
H1:p≠0.7

H0:μ≤0.7
H1:μ>0.7

H0:p≤0.7
H1:p>0.7

H0:μ=0.7
H1:μ≠0.7

H0:p≥0.7
H1:p<0.7
Answer #1

H0: p 0.7

Ha: p > 0.7

Test statistics

z = ( - p ) / sqrt [ p ( 1 - p) / n ]

= ( 0.71 - 0.70) / sqrt [ 0.7 * ( 1 - 0.7) / 200 ]

= 0.31

p-value = P(Z > z)

= P(Z > 0.31)

= 0.3783 (From Z table)

Since p-value > 0.05 , fail to reject H0